# Entangled in Representations: Mechanistic Investigation of Cultural Biases in Large Language Models

### v1

@@ 4.2 Cultural Flattening Score
LLMs trained on imbalanced cultural resources are likely to represent less-documented cultures through the cultural knowledge of more dominant ones, potentially leading to overgeneralization. To quantify this phenomenon, we introduce a Cultural Flattening score (CF score) that measures the extent to which one country’s learned representation has been homogenized to resemble another’s.
A CF score is asymmetric and calculated for a pair of countries, target country $y_{t}$ and source country $y_{s}$. To calculate the CF score, we first compute a cultural knowledge signature for each country. A cultural knowledge signature is a collection of cultural knowledge decoded from Culturescope (§4.1). Let $\mathcal{CK}$ denote the set of cultural knowledge and $\mathcal{Y}$ the set of all countries. For a country $y\in\mathcal{Y}$, let $S_{y}(ck)$ represent the set of activation scores obtained by Eq. 3 for knowledge $ck\in\mathcal{CK}_{y}$ across all instances from country $y$. We define the unnormalized knowledge signature for country $y$ as:
$\tilde{\sigma}_{y}(ck)=\begin{cases}\bar{s}_{y,ck}\cdot\log(1+|S_{y}(ck)|)&\text{if }S_{y}(ck)\neq\emptyset\\ 0&\text{otherwise}\end{cases}$ (4)
where $\bar{s}_{y,ck}=\frac{1}{|S_{y}(ck)|}\sum_{s\in S_{y}(ck)}s$ is the average activation score for knowledge $ck$ in country $y$, and $|S_{y}(ck)|$ denotes the frequency of knowledge $ck$ appearing in country $y$’s data.
The normalized cultural knowledge signature is then computed as:
$\sigma_{y}(ck)=\frac{\tilde{\sigma}_{y}(ck)}{\sum_{ck^{\prime}\in\mathcal{CK}}\tilde{\sigma}_{y}(ck^{\prime})}$ (5)
We define the set of universal concepts as those appearing in every country: $\mathcal{U}=\{ck\in\mathcal{CK}:\forall y\in\mathcal{Y},S_{y}(ck)\neq\emptyset\}$. The CF score from target country $y_{t}$ to source country $y_{s}$ is:
$F(y_{t}\rightarrow y_{s})=\sum_{ck\in\mathcal{CK}\setminus\mathcal{U}}\sigma_{y_{t}}(ck)\cdot\mathbb{1}[S_{y_{s}}(ck)\neq\emptyset]$ (6)
where $\mathbb{1}[\cdot]$ is the indicator function. This formulation ensures that $F(y_{t}\rightarrow y_{s})\in[0,1]$, with higher values indicating that a larger fraction of $y_{t}$’s culturally distinctive knowledge is shared with $y_{s}$.
This asymmetric nature of scores directs cultural influence: a high score $F(y_{t}\rightarrow y_{s})$ indicates that country $y_{t}$’s representation substantially overlaps with country $y_{s}$’s, suggesting that the model may have learned to represent $y_{t}$ through the $y_{s}$’s cultural knowledge.
@@ 5 Tracing LLMs’ Internal Mechanisms for Cultural Knowledge Usage
Leveraging MCQs with hard negatives (§3.2.3), we propose to investigate how LLMs internalize Western-dominance bias and cultural flattening via the attention map Yuksekgonul et al. (2024). This setup facilitates an analysis of whether cultural biases are also reflected in the attention mechanism, as in extrinsic evaluation, tracing the internal mechanisms for the emergence of cultural biases.
@@ 5.1 Attention Contribution Score
To examine how LLMs internally process Western-dominance bias and cultural flattening, we analyze how attention patterns are directed toward each option $c$ in the input $T$ using the MCQs with hard negatives which contains option list simulating the bias (§3.2.3). Following Yuksekgonul et al. (2024)’s work that highlights the final input token as a meaningful anchor point for attention analysis, we track attention from this final input token $t_{s}$ to tokens that correspond to options in an option list $C$.
For a tokenized input text $T$ that contains a culture-specific question $q$ and a curated option list, we compute the attention contribution $a_{t_{c},t_{s}}$ from the final input token $t_{s}$ to the token $t_{c}$ corresponding to an option $c$, using the following procedure. The operation involves four projection matrices $W^{L}_{Q},W^{L}_{K},W^{L}_{V},W^{L}_{O}\in\mathbb{R}^{d\times d}$ that correspond to the ‘query’, ‘key’, ‘value’, and ‘output’ projections in the attention block of the layer. Each of these is split into multiple heads, where $W^{l,h}_{Q},W^{l,h}_{K},W^{l,h}_{V}\in\mathbb{R}^{d\times d_{h}}$ and $W^{l,h}_{O}\in\mathbb{R}^{d_{h}\times d}$ denote the matrices for each head $h$. $H$ is the total number of heads, $d_{h}$ is the dimensionality for each head. Embeddings are split into equal parts such that $d_{H}=\frac{d}{H}$. The attention weight matrix $A^{l,h}$, calculated as in Eq. 7, is comprised of the attention weight values computed by the attention head $h$ at layer $l$ over an input $T$ containing $S$ number of tokens, where Softmax is taken row-wise. The layer-wise attention contribution score, $a_{t_{c},t_{s}}^{l}$, is defined as in Eq. 8, where $A^{l,h}_{t_{c},t_{s}}$ is the attention weight of the token $t_{c}$ and $t_{s}$ from the matrix $A^{l,h}$.
$A^{l,h}=\mathrm{Softmax}(\frac{(X^{l-1}W^{l,h}_{Q})(X^{l-1}W^{l,h}_{K})^{T}}{\sqrt{d_{h}/H}})$ (7)
$a_{t_{c},t_{s}}^{l}=\sum_{h=1}^{H}A^{l,h}_{t_{c},t_{s}}(x^{l-1}_{t_{c}}W^{l,h}_{V})W^{l,h}_{O}$ (8)
The final attention contribution score, $a_{t_{c}}$, is computed by averaging the layer-wise attention contribution scores $a_{t_{c},t_{s}}^{l}$ across all layers in the LLM, where $t_{s}$ is the final input token.
@@ 6 Experimental Results
BLEnD
@@
Baseline 0.7126 0.7019 0.6675
Cultural Prompting 0.6799 0.7282 0.7148
CANDLE 0.5640 0.6955 0.6473
Culturescope 0.6519 0.7169 0.6659
Table 1: We present the QA performances of three different LLMs with different inputs on BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We highlight the best performing method in bold, and the second-best in italics.
6.1 Open-Ended QA Performances
To ensure its relevance to the input, we prepend the comma-separated cultural knowledge $CK$ to the input text $T$. The input example can be found in Appendix B. Given the augmented input $\mathcal{CK}_{i};T_{i}]$, an LLM is asked to generate an output $O$. To evaluate an LLM’s answer $O$, we perform exact-match evaluation, whether $O$ contains a gold answer $a$. We compare our method to three different input schemes, which are designed to enhance LLMs’ cultural understanding ability. Table 1 shows the accuracy from different input schemes.
Baseline is an input without an explicit instruction and additional cultural knowledge. Cultural Prompting Li et al. (2024); Cheng et al. (2023) is an input with an additional instruction (e.g. “The following question is about country yy”) designed to guide an LLM with an explicit country name. CANDLE Nguyen et al. (2023b) is a comprehensive cultural commonsense knowledge base consisting of triples (country, topic, assertion) spanning 196 countries. It provides cultural concepts derived from assertions, which we compare to the cultural knowledge decoded from Culturescope.
Since topics in CANDLE and those in the datasets are not aligned, we sample 20 concepts for each country. Due to the sampling, we perform the inference with three different random seeds and report the average accuracy. Experimental results show that inputs augmented with Culturescope perform better than those with CANDLE concepts. The accuracy with Culturescope is best on BLEnD across all models, on par on CAMeL-2 dataset with the baseline. This confirms that Culturescope reveals a valid cultural knowledge space, which is highly relevant to the given input.
Figure 3: We present the results from the CF score on BLEnD with Llama-3.1. A connection from $y_{t}$ to $y_{s}$ means that the fraction of $y_{t}$’s distinctive knowledge results from the cultural flattening with $y_{s}$.
6.2 Results with CF Score
Figure 3 illustrates the country pairs that exhibit cultural flattening, as measured by our CF score. To highlight only the most significant connections, we apply a threshold to exclude pairs with CF scores lower than the average of all country pairs. Notably, Assam and Indonesia are omitted from the visualization due to their relatively low scores. Overall, the results suggest that countries within the same regional groups tend to share a common cultural knowledge space with bidirectional connections. Additionally, Iran, the United States, and China appear prominently as source countries ($y_{s}$), indicating their broader influence on the connected target countries ($y_{t}$). Further results using CAMeL-2 and other models are available in Appendix C.
@@ 6.2 Results with CF Score
Acc % Biased % Others Refusal Acc % Biased % Others Refusal
Llama-3.1 avg. 0.43 0.19 0.18 0.02 0.43 0.20 0.18 0.02
mid 0.44 0.19 0.18 0.02 0.45 0.21 0.16 0.02
low 0.39 0.21 0.19 0.02 0.37 0.19 0.21 0.03
aya-expanse avg. 0.38 0.18 0.16 0.11 0.35 0.19 0.17 0.13
mid 0.40 0.17 0.16 0.11 0.37 0.20 0.15 0.13
low 0.31 0.20 0.19 0.12 0.30 0.19 0.19 0.13
Qwen2.5 avg. 0.44 0.20 0.16 0.05 0.44 0.20 0.16 0.05
mid 0.44 0.19 0.17 0.04 0.47 0.19 0.15 0.04
low 0.40 0.21 0.17 0.06 0.36 0.21 0.18 0.06
Table 2: Model outputs result from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region dataset (§3.2.3), evaluated using four metrics. Results are aggregated by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low).
6.3 Performances on MCQ with Hard Negatives
Table 2 shows the model’s final output results from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region (§3.2.3). We aggregate the results by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low). This reflects LLM’s output-level preference for hard negative options over other options, which serves to reveal how internal biases within the LLM can affect its final prediction. We present the percentage of instances where the model chooses a gold answer, denoted as Accuracy (Acc). The metric labeled as “% Biased” indicates the proportion of instances where the model chooses a hard negative option, which represents a targeted bias. “% Others” represents the proportion of instances where the model chooses one of the remaining random options. Since there are two random options present in the option list, we divide the proportion of choosing random options by two for a fair comparison. We also report the proportion of instances as ‘Refusal’ where LLMs avoid answering.
In most cases, LLMs prefer a hard negative option to other options when they are generating wrong answers with a higher % Biased than % Others. By breaking down the results by resource levels Joshi et al. (2020), we observe that the accuracy decreases for low-resource target questions compared to the average. We also find that in BLEnD-resource, the proportion for selecting a hard negative option from high-resource cultures increases, suggesting the presence of Western-dominance bias. However, this trend does not necessarily extend to BLEnD-region, where the selection rate of hard negatives does not show a similar increase. This indicates that low-resource cultures may be less susceptible to cultural flattening, likely due to the limited cultural knowledge available in the model.
(a) BLEnD-Resource
(b) BLEnD-Region
Figure 4: We present a heatmap visualization of attention contribution scores (z-score normalized) for the Llama-3.1 model on incorrect predictions. Groups on x-axis represent the group of the country which the gold answer is from. Groups on y-axis represent the group of the country which the chosen option is from. For example, in Figure 4 (a), ‘Low (x-axis)’-‘High (y-axis)’ pair with the score of 0.44 shows the averaged attention contribution score on option tokens from ‘High’ resource group when the gold answer is from ‘Low’ resource group. In BLEnD-Resource, we observe high attention contribution scores to tokens from high-resource cultures. BLEnD-Region also shows similar trends, higher scores to Europe and North America, regional group consisting of high-resource countries.
6.4 Attention Contribution Score Analysis
Figure 4 presents the average of attention contribution scores on option token positions, assigned by Llama-3.1 when the model makes incorrect predictions. We separate the analysis between correct predictions and incorrect predictions, as we are particularly interested in LLMs’ internal patterns when they are making biased predictions. Details for the aggregation method and results from aya-expanse and Qwen2.5 are shown in Appendix D.
In both BLEnD-Resource and BLEnD-Region, groups on x-axis represent groups of the question’s target country—resource level (High, Mid, Low) in BLEnD-Resource and region (South Asia [S-AS], East Asia [E-AS], West Asia [W-AS], Europe [EUR], America [AME], Africa [AFR]) in BLEnD-Region — while groups on y-axis indicate groups of the country which the chosen option is from. In Figure 4 (a), visualizing results of BLEnD-resource, we find that attention contributions from the last input token to incorrect high-resource country options are higher than incorrect mid- and low-resource countries options, especially for low-resource target questions. Figure 4 (b), visualizing results of BLEnD-region, shows similar trends to Figure 4 (a). Llama-3.1 allocates higher attention contributions to European and North American countries options compared to other region groups options. This trend is also shown in other two models as in Figure 7.
The analysis of attention contributions demonstrates that high-resource bias and Western-dominance bias is highly internalized within LLMs’ representations. This contradicts the evaluation on LLMs’ performance in Table 2, where LLMs are equally exhibiting both biases. This further aligns with findings from societal biases where they find intrinsic bias metrics and extrinsic bias metrics do not always correlate Goldfarb-Tarrant et al. (2021).
7 Discussion
7.1 Cultural Flattening within LLMs’ Inner Representations
With our Culturescope method, we can now probe the cultural knowledge encoded within the internal representations of LLMs. As shown in Figure 3, which visualizes the cultural flattening direction between cultures, we find unidirectional connections that have Iran and the United States as $y_{s}$. This unidirectional connection implies that models may have learned to represent less documented cultures, such as Ethiopia and Algeria, through those high-resource cultures. Our experiments using hard negative options align with previous works, which find that LLMs sometimes respond with answers aligned with culturally similar or geographically proximate regions Cao et al. (2023); Tao et al. (2024). We further attribute the models’ tendency to favor culturally adjacent answers to the unidirectional connections found by the proposed Culturescope. These findings underscore the need for methods that can disentangle culturally entangled representations, particularly among similar cultures, to enhance the accuracy and cultural appropriateness of LLM outputs.
7.2 LLMs’ Performance on Low-Resource Cultures
Previous studies Azime et al. (2025); Myung et al. (2024); Li et al. (2024) have shown that LLMs often struggle to utilize knowledge relevant to low-resource cultures. Our MCQ results with hard-negative options (Table 2; §6.3) are consistent with these findings, reinforcing the narrative that LLMs underperform on cultural tasks involving low-resource regions. Additionally, the relatively low cultural flattening (CF) scores for low-resource cultures (§6.2) suggest that these models are less prone to generate culturally flattened outputs due to the limited cultural knowledge encoded in their parameters. This is further supported by the reduced ratio of biased answers with region-type hard negatives for low-resource cultures, as shown in Table 2. These findings indicate that LLMs exhibit weaker cultural biases toward low-resource cultures, not because of improved fairness, but due to a lack of cultural representation. Consequently, improving performance on low-resource cultures may require a different approach — one that prioritizes knowledge acquisition over bias mitigation.
8 Conclusion
In this work, we investigate the complex and often biased ways LLMs process cultural knowledge. We introduce Culturescope, a method that leverages activation patching to probe the internal mechanisms of LLMs, allowing us to analyse the cultural knowledge encoded within their layers. In addition, we quantify the phenomenon of ‘cultural flattening’, where LLMs represent less-documented cultures through the concepts of more dominant or geographically close ones, thereby erasing cultural nuances. We conduct our research across three distinct models using culturally-grounded QA datasets, moving beyond isolated extrinsic evaluations to examine the interaction between cultural knowledge within the models’ parametric space. Our analysis reveals that models over-attend Western-centric tokens internally, indicating internalized Western-dominance bias. The results for low-resource cultures suggest that they may not be as susceptible to cultural biases as other cultures. Our findings suggest that future work should develop a tailored approach that considers the impact of bias and resource levels to improve LLMs’ cultural understanding.
Limitations
While our study provides new insights into cultural knowledge in LLMs, there exists limitations. Although a growing number of benchmarks aim to evaluate cultural knowledge in LLMs, few are suitable for our evaluation setup. To meaningfully compare the probed cultural knowledge across cultures, the datasets must maintain a consistent QA format across different cultures. In addition, to ensure that the evaluation captures how LLMs represent culture on a global scale, it is essential to include a geographically diverse set of cultures. However, we emphasize that our proposed method is model- and task-agnostic, and can be applied to any dataset that meets these requirements.
Due to computational constraints, we are reporting results with  8B models, unable to conduct experiments on larger-scale models. In addition, our analysis does not cover all countries globally. This limitation is not due to methodological oversight but rather the lack of publicly available datasets that include culturally grounded QA data for many regions. As such, our findings are necessarily constrained by the scope of existing resources. Finally, our MCQ with hard negatives (§3.2.3) involves a degree of random sampling. While this introduces some variability, we consider it a reasonable trade-off given the prohibitive cost of exhaustively evaluating all possible negative combinations. We mitigate this by ensuring consistency across runs and focusing on aggregate trends rather than individual instances.
Acknowledgements
[…]
Appendix A Dataset Details
@@
Extractive QA 1862 Locations, Beverage, Food, Sports

### v2

@@ 4.2 Cultural Flattening Score
To quantify cultural flattening, where a culture’s distinctive characteristics are misrepresented by those of a dominant culture, we introduce the Cultural Flattening score (CF score). A CF score is asymmetric and is calculated for an ordered pair of cultures, a target culture $y_{t}$ and a source culture $y_{s}$.
For the CF score, we first compute a chi-square contribution for each cultural knowledge decoded by Culturescope (Sec. 4.1). This chi-square contribution quantifies how strongly a given piece of knowledge is activated for a specific culture relative to an independence assumption across all cultures, thereby serving as a measure of cultural specificity. This formulation highlights the impact of culturally distinctive knowledge rather than shared cultural knowledge arising from cultural assimilation.
For knowledge $k$ from a culture $y\in\mathcal{Y}$, the expected count under the independence assumption, $\mathbb{E}_{k,y}$, is defined as:
$\mathbb{E}_{k,y}=\frac{N_{y}\cdot\sum_{y}g_{k,y}}{N_{\mathcal{Y}}}$ (4)
where $g_{k,y}$ is the frequency of $k$ in the culture $y$, and $N$ denotes the number of knowledge generated by $y$ or across all cultures $\mathcal{Y}$. The chi-square contribution of cultural knowledge $\mathcal{X}_{k,y}$ is then computed as:
$\mathcal{X}_{k,y}=\frac{(max(0,g_{k,y}-\mathbb{E}_{k,y}))^{2}}{\mathbb{E}_{k,y}}$ (5)
We only consider cases where the residuals are positive, in order to remove knowledge that is underrepresented by the culture. The higher the $\mathcal{X}_{k,y}$, the more distincitve $k$ is. The final CF score from the target culture $y_{t}$ to the source culture $y_{s}$ is:
$F(y_{t}\rightarrow y_{s})=\sum_{k\in\mathcal{K}}\mathcal{X}_{k,y_{t}}\cdot\mathbb{1}[\mathcal{X}_{k,y_{s}}\neq 0]$ (6)
where $\mathbb{1}[\cdot]$ is the indicator function. By summing chi-square contribution scores over overlapping knowledge, the CF score captures the total amount of culturally distinctive knowledge from the target culture that is also present in the source culture. Higher values indicate that a large fraction of the target culture’s distinctive concepts are represented in the source culture.
@@ 5 Tracing LLMs’ Internal Mechanisms for Cultural Knowledge Usage
Leveraging MCQs with hard negatives (Sec. 3.2.3), we propose to investigate how LLMs internalize Western-dominance bias and cultural flattening via the attention map Yuksekgonul et al. (2024). This setup facilitates an analysis of whether cultural biases are also reflected in the attention mechanism, as in extrinsic evaluation, tracing the internal mechanisms for the emergence of cultural biases.
@@ 5.1 Attention Contribution Score
To examine how LLMs internally process Western-dominance bias and cultural flattening, we analyze how attention patterns are directed toward each option $c$ in the input $T$ using the MCQs with hard negatives (Sec. 3.2.3). Following Yuksekgonul et al. (2024)’s work that highlights the final input token as a meaningful anchor point for attention analysis, we track attention from this final input token $t_{s}$ to tokens that correspond to $c$ in an option list $C$. To avoid biases from MCQ Wang et al. (2024b), we provide options without option letters.
For a tokenized input text $T$ that contains a culture-specific question $q$ and a curated option list, we compute the attention contribution $a_{t_{c},t_{s}}$ from the final input token $t_{s}$ to the token $t_{c}$ corresponding to an option $c$, using the attention weight of the token $t_{c}$ and $t_{s}$. As in Eq. 7, we obtain a layer-wise attention contribution score $a_{t_{c},t_{s}}^{l}$ on layer $l$, where $ATTN_{t_{c},t_{s}}$ denotes attention operation between the token $t_{c}$ and $t_{s}$ with a hidden representation $x^{l-1}_{t_{c}}$ from the previous layer.
$a_{t_{c},t_{s}}^{l}=ATTN_{t_{c},t_{s}}(x^{l-1}_{t_{c}})$ (7)
The final attention contribution score, $a_{t_{c}}$, is computed by averaging the layer-wise scores $a_{t_{c},t_{s}}^{l}$ across all layers in the LLM.
Figure 3: We present the results from the CF score on BLEnD with Llama-3.1 in English. Percentages on the right reflect relative knowledge coverage and are defined as the ratio between the number of cultural knowledge generated for a given culture and the total number of generated concepts across all cultures.
@@ 6 Experimental Results
6.1 Intrinsic Cultural Bias Evaluation
6.1.1 CF Score Results
We present CF score results using a Sankey diagram, where the flow direction indicates target-to-source relations and flow width is proportional to the CF score. Figure 3 shows cultural pairs with high CF scores, where strand width reflects the extent to which culturally distinctive knowledge from a target culture $y_{t}$ is generalized to a source culture $y_{s}$ in the model’s internal representations.
Despite their large number of generated concepts, the US and the UK exhibit relatively weak incoming connections as source countries, indicating that the CF score successfully downweights overlap driven by broadly shared cultural knowledge. The result reveals asymmetric connections, often aligned with geographic or cultural proximity, suggesting uneven generalization of culturally distinctive knowledge (China $\rightarrow$ S. Korea, Iran $\rightarrow$ Azerbaijan). In addition, cultures with lower concept coverage, such as Assam, Azerbaijan, and Ethiopia, consistently show narrow connections. This pattern of asymmetry is also shown in other languages. Additional results using CAMeL-2 and other models, covering English, Arabic, and Spanish, are provided in Appendix E.
(a) BLEnD-Resource
(b) BLEnD-Region
Figure 4: We present a heatmap of attention contribution scores (z-score normalized) for incorrect predictions of Llama-3.1. The x-axis groups correspond to the culture of the chosen option, while the y-axis groups correspond to the culture of the gold answer. For example, in Figure 4 (a), the High-Low cell (0.44) indicates the average attention contribution to option tokens from the High resource group when the gold answer belongs to the Low resource group.
BLEnD-Resource BLEnD-Region
Acc % Biased % Others Refusal Acc % Biased % Others Refusal
Llama-3.1 avg. 0.43 0.19** 0.18 0.02 0.43 0.20** 0.18 0.02
mid 0.44 0.19 0.18 0.02 0.45 0.21** 0.16 0.02
low 0.39 0.21 0.19 0.02 0.37 0.19 0.21 0.03
aya-expanse avg. 0.38 0.18** 0.16 0.11 0.35 0.19** 0.17 0.13
mid 0.40 0.17 0.16 0.11 0.37 0.20** 0.15 0.13
low 0.31 0.20 0.19 0.12 0.30 0.19 0.19 0.13
Qwen2.5 avg. 0.44 0.20** 0.16 0.05 0.44 0.20** 0.16 0.05
mid 0.44 0.19 0.17 0.04 0.47 0.19** 0.15 0.04
low 0.40 0.21** 0.17 0.06 0.36 0.21** 0.18 0.06
Table 1: Model outputs result from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region dataset (Sec. 3.2.3), evaluated using four metrics. Results are aggregated by question’s target culture type: all averaged (avg.), mid-resource (mid), and low-resource (low). ** denotes statistically significant difference between % Biased and % Others.
6.1.2 Attention Contribution Score Results
Figure 4 presents the average of attention contribution scores on option token positions, assigned by Llama-3.1 when the model makes incorrect predictions. We separate the analysis between correct predictions and incorrect predictions, as we are particularly interested in LLMs’ internal patterns when they are making biased predictions. Details for the aggregation method and results from aya-expanse and Qwen2.5 are shown in Appendix F.
In BLEnD-Resource and BLEnD-Region, the x-axis groups correspond to the groups of the chosen option, while the y-axis groups represent the target culture of the question—resource level (High, Mid, Low) in BLEnD-Resource and region (South Asia [S-AS], East Asia [E-AS], West Asia [W-AS], Europe [EUR], America [AME], Africa [AFR]) in BLEnD-Region. From Figure 4 (a) and (b), attention contributions from the last input token to incorrect options reflect the induced biases, Western-dominance bias in BLEnD-Resource and cultural flattening in BLEnD-Region. This trend is also shown in two additional models (Figure 8). Student’s t-tests indicate that the average differences in attention scores are statistically significant. Details are provided in Appendix D.2.
6.2 Extrinsic Cultural Bias Evaluation
6.2.1 Performances on MCQ with Hard Negatives
Table 1 shows the extrinsic evaluation results from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region (Sec. 3.2.3). We aggregate the results by question’s target culture type: all averaged (avg.), mid-resource (mid), and low-resource (low). We present the percentage of instances where the model chooses a gold answer, denoted as Accuracy (Acc). The metric labeled as “% Biased” indicates the proportion of instances where the model chooses a hard negative option, which represents a targeted bias. “% Others” represents the proportion of instances where the model chooses one of the remaining random options. Since there are two random options present in the option list, we divide the proportion of choosing random options by two for a fair comparison. We also report the proportion of instances as ‘Refusal’ where LLMs avoid answering.
We observe that the accuracy decreases for low-resource target questions compared to the average, which is statistically significant (Table 7). We further statistically test whether the lower accuracy in low-resource cultures stems from a preference for biased options over others. However, while the differences between % Biased and % Others are statistically significant in avg. and mid-resource cultures, the differences in low-resource cultures are not statistically significant. Details for the statistical tests are described in Appendix D.1.
7 Evaluation of Culturescope
Considering that Culturescope elicits cultural knowledge from all layers without pinpointing any specific model component, we design two validation approaches including relevance evaluation and irrelevant patching. First, for the relevance evaluation, we compare our method to three different input schemes, which are designed to enhance LLMs’ cultural understanding ability. Specifically, For Culturescope, we prepend the comma-separated cultural knowledge $K$ to the input text $T$ and evaluate an LLM’s answer $O$ by exact-match. Table 2 shows the accuracy from different input schemes on BLEnD. Results on CAMeL-2 are in Table 9.
Baseline is an input without an explicit instruction and additional cultural knowledge. Cultural Prompting Li et al. (2024); Cheng et al. (2023) is an input with an additional instruction (e.g. “The following question is about culture yy”) designed to guide an LLM with an explicit culture name. CANDLE Nguyen et al. (2023b) is a comprehensive cultural commonsense knowledge base which provides cultural concepts derived from assertions. The input example can be found in Appendix B.
Since topics in CANDLE and those in the datasets are not aligned, we sample 20 concepts for each culture. Due to the sampling, we perform the inference with three different random seeds and report the average accuracy. Experimental results show that inputs augmented with Culturescope perform better than those with CANDLE concepts. The accuracy with Culturescope is best on BLEnD across all models, on par on CAMeL-2 dataset with the baseline. This confirms that Culturescope reveals a valid cultural knowledge space, which is highly relevant to the given input.
For the irrelevant patching test, instead of patching representative hidden representations, we patch randomly sampled Gausian noise vectors and decode knowledge using the same inspection prompt. We report the average semantic similarity score of the generated knowledge computed during the filtering stage. Higher similarity scores for Culturescope validate that the generated cultural knowledge is not an artifact induced by the experimental setup (Table 2; Table 8). In addition, we provide the average number of generated knowledge per layer in Appendix C.
Relevance evaluation
@@ 7 Evaluation of Culturescope
English 0.325 $\pm$ 0.01 0.349 $\pm$ 0.1 0.344 $\pm$ 0.11
Irrelevant 0.149 $\pm$ 0.09 0.127 $\pm$ 0.07 0.141 $\pm$ 0.09
Table 2: We present the evaluation of Culturescope on BLEnD Myung et al. (2024) in English. For the relevance evaluation results, we highlight the best-performing method in bold and the second-best in underline. For the irrelevant patching results, we report the mean and standard deviation.
8 Discussion
Our proposed methods for evaluating intrinsic cultural biases in LLMs reveal that models internalize two cultural biases: Wester-dominance bias and cultural flattening. As shown in Figure 3, we observe asymmetric connections among geographically or culturally proximate cultures. These asymmetries indicate uneven generalization of culturally distinctive knowledge, consistent with potential cultural flattening effects. Attention-based analyses further reveal the presence of Western-dominance bias that is not captured by CF scores as the CF score explicitly downweights globally shared knowledge (Section 6.1.2; Figure 4). This highlights that the CF score and attention contribution scores capture different but complementary aspects of cultural bias in model representations.
We also observe narrow connections for cultures with low coverage, such as Assam, Azerbaijan, and Ethiopia (Section 6.1.1). Rather than indicating cultural erasure through dominance by other cultures, this pattern likely reflects limited representational coverage of culturally distinctive knowledge of LLMs. In this sense, low-resource cultures may appear less susceptible to cultural flattening. This interpretation aligns with Li et al. (2025), who report weaker memorization of symbols associated with low-resource cultures in LLMs.
Further evidence for this explanation comes from downstream behaviour: models exhibit a reduced preference for biased options involving low-resource cultures (Table 1), with the differences being statistically significant (Table 7). Taken together, these findings suggest that weaker cultural biases toward low-resource cultures do not stem from improved fairness, but rather from insufficient cultural representation. Consequently, improving model performance for low-resource cultures may require approaches that prioritize knowledge acquisition and representational coverage, rather than bias mitigation alone.
9 Conclusion
In this work, we study how cultural biases arise from the internal representations of LLMs, by investigating LLMs’ internal representations. We introduce Culturescope, a method that leverages activation patching to analyse the cultural knowledge encoded within their layers. In addition, we quantify the phenomenon of ‘cultural flattening’, where LLMs represent less-documented cultures through the knowledge of more dominant or geographically close ones, thereby erasing cultural nuances. Our findings suggest that future work should develop a tailored approach that considers the impact of bias and resource levels to improve LLMs’ cultural understanding.
Limitations
Although a growing number of benchmarks aim to evaluate cultural knowledge in LLMs, few are suitable for our evaluation setup. To meaningfully compare the probed cultural knowledge across cultures, the datasets must maintain a consistent QA format across different cultures. Thus, our methods and analyses inherit the limitations from the benchmarks, such as, lack of fine-grained set of cultures, and 1-to-1 mapping of a culture and a country. However, we emphasize that our proposed method is model- and task-agnostic, and can be applied to any dataset that meets these requirements.
Due to computational constraints, we are reporting results with  8B models, unable to conduct experiments on larger-scale models. Our definition of low-resource cultures relies on the approximation of the most-spoken language from the previous work Joshi et al. (2020). Finally, our MCQ with hard negatives (Sec. 3.2.3) involves a degree of random sampling. While this introduces some variability, we consider it a reasonable trade-off given the prohibitive cost of exhaustively evaluating all possible negative combinations. We mitigate this by ensuring consistency across runs and focusing on aggregate trends rather than individual instances.
Appendix A Dataset Details
In our experiments, we utilize BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). BLEnD comprises 500 short-answer question-answer pairs for each culture, where the answers vary depending on the cultural or regional context. To reduce computational cost, we exclude North Korea and West Java, resulting in a final selection of 14 cultural groups from BLEnD.
With CAMeL-2, an LLM is asked to extract an entity from a context collected from Arabic X/Twitter data according to the specified entity type in the input. We take the English partition and reduce the dataset to 14 countries to keep a similar culture distribution to the BLEnD dataset. Dataset details, including domains and countries covered by the datasets, can be found in Appendix A.
We provide their brief data statistics and characteristics in Table 6.
A.1 Culture Groups
As mentioned in Sec. 3.2, our work conducts experiments that focus on 14 countries classified in two dimensions. Specifically, with the taxonomy from Joshi et al. (2020) which classifies languages into six levels (0: very low-resource to 5: very high-resource), we simplify this into three groups: High (Level 5), Mid (Levels 3-4), and Low (Levels 0-2). We assign each culture a language resource level based on its most widely spoken language, provided by Wikipedia.
For the region dimension, we group countries into six regions based on continents. We split Asia into three subregions, which leaves us six regional groups: North America, Europe, Africa, West Asia, South Asia, and East Asia. We compare three groups based on the level of language resource Joshi et al. (2020) and six groups based on the continental region. Table 3 shows the culture entities that correspond to each group.
@@ A.1 Culture Groups
High 2548 Europe 789
North America 520
Mid 1425 East Asia 588
West Asia 516
Low 1077 South Asia 592
Africa 771
Sum 5050 Sum 3776
Table 4: Number of questions per group in Cultural MCQ with hard negatives
Option type Mean $\pm$ s.d. 95% CI
BLEnD-Region
Gold answer 269.71 $\pm$ 16.49 [260.2, 279.23]
Hard negative 269.71 $\pm$ 16.3 [260.3, 279.12]
Others 539.43 $\pm$ 38.34 [517.3, 561.56]
BLEnD-Resource
Gold answer 360.71 $\pm$ 11.4 [354.13, 367.30]
High 721.43 $\pm$ 31.55 [692.25, 750.61]
Mid 1683.33 $\pm$ 52.5 [1553.9, 1812.78]
Low 1262.5 $\pm$ 85.57 [1126.49, 1398.51]
Table 5: Descriptive statistics (mean, standard deviation (s.d.), and 95% Confidence Interval (CI)) of the number of selected culture groups for each option type in BLEnD-Resource and BLEnD-Region.
@@
Extractive QA 1,862 Locations, Beverage, Food, Sports